OpenStudy (anonymous):

Find t9 for the geometric sequence: 189, 63, 21, 7, … Give your answer as a decimal to the nearest hundreth

OpenStudy (anonymous):

r=3 a=189

OpenStudy (dumbcow):

r= 1/3, series is decresing

OpenStudy (anonymous):

equation t(9)=189(1/3)*(9-1)

OpenStudy (dumbcow):

\[a_{n} = a_{1}*r^{n-1}\]

OpenStudy (angela210793):

189/63=3 t9==t1*r^(n-1)=189*1/3^8

OpenStudy (anonymous):

is it 0.03

OpenStudy (anonymous):

is 0.03 right